# FetchBench: A Simulation Benchmark for Robot Fetching

### v1

@@ Abstract
Fetching, which includes approaching, grasping, and retrieving, is a critical challenge for robot manipulation tasks. Existing methods primarily focus on table-top scenarios, which do not adequately capture the complexities of environments where both grasping and planning are essential. To address this gap, we propose a new benchmark FetchBench, featuring diverse procedural scenes that integrate both grasping and motion planning challenges. Additionally, FetchBench includes a data generation pipeline that collects successful fetch trajectories for use in imitation learning methods. We implement multiple baselines from the traditional sense-plan-act pipeline to end-to-end behavior models. Our empirical analysis reveals that these methods achieve a maximum success rate of only 20%, indicating substantial room for improvement. Additionally, we identify key bottlenecks within the sense-plan-act pipeline and make recommendations based on the systematic analysis.
@@ 1 Introduction
Based on the benchmark, we implement and evaluate several baselines for the task. In particular, the common pipeline used in previous works [1, 11, 5] is to use grasp pose prediction models to propose candidate grasp poses and then use motion planning [12] algorithms to search for the collision-free motion to reach the target location. We find in Section 6 that this pipeline is no longer reliable when fetching from novel and complex environments, achieving only 13% success rate. We analyze the failure cases by placing oracle components within the pipeline and point out several challenges that remain on existing motion planning algorithms and grasp pose prediction models (Section 6.2).
@@ 2 Related Works
Manipulation Benchmark in Simulators. Benchmarking robotic systems and algorithms in simulation has the advantage of high efficiency, low cost, and reproducibility. Compared to existing benchmarks [14, 17, 31, 22, 18, 24, 15, 16, 27, 29] for robot manipulation (Table 1), our benchmark is unique in the following aspects: First, we specialize in evaluating robot fetching of unknown objects from novel environments, while other benchmarks either focus on general-purpose robot learning algorithms, e.g. imitation learning [22, 18, 17, 27], reinforcement learning [18, 14, 24, 16, 17, 29], or on long-horizon embodied AI challenge by simplifying the grasping process [15, 31]. Second, our benchmark provides abundant procedural environments for training and evaluation, with a total of 5.5k objects for grasping.
Other benchmarks are limited by the number of artist-designed environments and the number of objects to be grasped.
@@ 4.1 Procedural Task Generation
We create 13 types of procedural scenes made with objects such as shelves, cabinets, drawers, dressers, baskets, etc. Each scene type combines several different procedural assets from Infinigen [40, 10] [^1] We received early access to the code from the authors.. Our scene generation procedure has a total of 310+ independent parameters that can be randomly sampled to create diverse scene instances. Next, we automatically annotate support surfaces (e.g., table-tops, shelf boards) in procedural scenes to spawn the objects. In particular, there are four categories of support surfaces: on-table, on-shelf, in-drawer, and in-basket. These labels allow us to evaluate methods based on scene category. For more details on procedural scene generation, please refer to Appendix A.2.
@@
We note that some task instances might be infeasible to solve based on stability at initialization, so we filter out instances where (1) an object’s initial pose is not static (i.e. the object is falling off after initialization), (2) a collision-free IK solution does not exist for all annotated grasp poses, and (3) the target object is near absent in the input point cloud.
In all, we generate a total of 6k tasks for testing in our benchmark: 1526 on-table cases, 2724 on-shelf cases, 891 in-basket cases, and 859 in-drawer cases. Appendix A.3 shows more examples.
@@ 5 Baselines
ContactGraspNet-Cabinet. Kumaar et al. [9] proposes a neural collision checker for partially observed objects and scenes. We re-implement the same pipeline as above, but use MPPI [44] iterations with Cabinet’s neural collision checking and waypoints proposals as the motion planning module.
End-to-End Behavior Cloning. We train an end-to-end policy to grasp the object from the cluster with expert demos. In our architecture, we first compute the embeddings for the segmented point cloud using PointNet++ [45] encoders and concatenate them with joint states and end-effector embeddings. These embeddings are then passed to a backbone to output the delta joint movements at 10 Hz. We use two variants for the backbone: an MLP [19] that takes the current step embeddings, and a transformer [46] that takes in a history of embeddings. We name these E2EImit-MLP and E2EImit-Transformer, respectively.
@@ 6.1 Baseline Comparisons
Next, we observe that the end-to-end behavior cloned policies have <<<2% success rate. Specifically, we find that the grasping phase presents significant challenges for an end-to-end trained model, as the model must learn both collision-free planning and generalize to grasp various object shapes. Thus, our dataset and benchmark offer a challenging testbed for researching imitation learning methods [37, 46] in object fetching, where the methods can be evaluated based on their ability to generalize to unknown objects and scenes.
@@
Table 2: Evaluation of baselines on our benchmark. We show the average success rate, the average computation time and trajectory length on the success cases.
Total Succ Computation Time (s) Traj Length (Rad)
@@
E2EImit-MLP 0.018 19.8 15.2
E2EImit-Transformer 0.012 18.3 17.9
CGN-CuRobo-Imit-MLP 0.189 16.7 6.6
CGN-CuRobo-Imit-Transformer 0.202 16.7 6.41
@@ 7 Real Robot Experiment
In total, we achieve a success rate of 21%, including 38% on on-table tasks, 13.3% on on-shelf tasks and 12.5% on in-basket tasks. The result suggests the challenge of fetching unknown objects from complex and cluttered scenes (e.g., shelves, baskets) in the real-world. Moreover, out of all the failure cases, 25.8 % comes out of low-quality grasp poses (the object slips away), 27.3 % comes out of motion planning failure, 23.5 % comes from unexpected collision with the scene, and the remaining comes from no grasp proposals. We refer to Appendix C for the experiment details, evaluations and videos.
@@ B.2 Imitation Learning
For our imitation learning models, we use PointNet++ [45] as the encoder on point cloud of the target object, the scene and the robot separately. Furthermore, we encode the proprioceptive states with a MLP encoder. We concatenate all embeddings, i.e., target object point cloud embedding, scene point cloud embedding, robot point cloud embedding and the proprioceptive state embedding. Then, we project it to a fixed size as the final embedding to feed into the policy network.
For the MLP variant, we use the embedding of current step and use a 3-layer MLP to output the delta joint movement. The output is parameterized as a Tanh Gaussian distribution. For the Transformer variant, we use the consecutive last 4 steps’ embeddings to feed into an Optimus policy head [46] and output a Tanh Gaussian mixture distribution for the joint movement.
Additionally, we downsample the input point cloud with farthest point sampling and the delta joint commands are re-scaled to fit the Tanh distribution. For E2EImit methods, we train a approach phase model and a retrieval phase model separately. The retrieval phase model is also used for the CGN-CuRobo-Imit methods. All models are trained on our demonstration dataset with Adam [47] optimizer. We select the best hyper-parameters by validating on held-out validation tasks.
Appendix C Real Robot Experiment
C.1 Experiment Setup
@@ C.2 Experiment Results
Appendix D Additional Related Works

### v2

@@ Abstract
Fetching, which includes approaching, grasping, and retrieving, is a critical challenge for robot manipulation tasks. Existing methods primarily focus on table-top scenarios, which do not adequately capture the complexities of environments where both grasping and planning are essential. To address this gap, we propose a new benchmark FetchBench, featuring diverse procedural scenes that integrate both grasping and motion planning challenges. Additionally, FetchBench includes a data generation pipeline that collects successful fetch trajectories for use in imitation learning methods. We implement multiple baselines from the traditional sense-plan-act pipeline to end-to-end behavior models. Our empirical analysis reveals that these methods achieve a maximum success rate of only 20%, indicating substantial room for improvement. Additionally, we identify key bottlenecks within the sense-plan-act pipeline and make recommendations based on the systematic analysis. The code for the benchmark is available at https://github.com/princeton-vl/FetchBench-CORL2024 .
@@ 1 Introduction
Based on the benchmark, we implement and evaluate several baselines for the task. In particular, the common pipeline used in previous works [1, 11, 5] is to use grasp pose prediction models to propose candidate grasp poses and then use motion planning [12] algorithms to search for the collision-free motion to reach the target location. We find in Section 6 that this pipeline is no longer reliable when fetching from novel, complex, and partially-observed environments, achieving only 13% success rate. We analyze the failure cases by placing oracle components within the pipeline and point out several challenges that remain on existing motion planning algorithms and grasp pose prediction models (Section 6.2).
@@ 2 Related Works
Manipulation Benchmark in Simulators. Benchmarking robotic systems and algorithms in simulation has the advantage of high efficiency, low cost, and reproducibility. Compared to existing benchmarks [14, 17, 31, 22, 18, 24, 15, 16, 27, 29, 21] for robot manipulation (Table 1), our benchmark is unique in the following aspects: First, we specialize in evaluating robot fetching of unknown objects from novel environments, while other benchmarks either focus on general-purpose robot learning algorithms, e.g. imitation learning [22, 18, 17, 27], reinforcement learning [18, 14, 24, 16, 17, 29], or on long-horizon embodied AI challenge by simplifying the grasping process [15, 31]. Second, our benchmark provides abundant procedural environments for training and evaluation, with a total of 5.5k objects for grasping. Other benchmarks are limited by the number of artist-designed environments and the number of objects to be grasped.
@@ 4.1 Procedural Task Generation
We create 13 types of procedural scenes made with objects such as shelves, cabinets, drawers, dressers, baskets, etc. Each scene type combines several different procedural assets from Infinigen [40, 10]. Our scene generation procedure has a total of 310+ independent parameters that can be randomly sampled to create diverse scene instances. Next, we automatically annotate support surfaces (e.g., table-tops, shelf boards) in procedural scenes to spawn the objects. In particular, there are four categories of support surfaces: on-table, on-shelf, in-drawer, and in-basket. These labels allow us to evaluate methods based on scene category. For more details on procedural scene generation, please refer to Appendix A.2.
@@
We note that some task instances might be infeasible to solve, so we filter out instances where (1) an object’s initial pose is not static (i.e. the object is falling off after initialization), (2) a collision-free IK solution does not exist for all annotated grasp poses, and (3) the target object is near absent in the input point cloud. In all, we generate a total of 6k tasks for testing in our benchmark: 1526 on-table cases, 2724 on-shelf cases, 891 in-basket cases, and 859 in-drawer cases. Appendix A.3 shows more examples.
@@ 5 Baselines
ContactGraspNet-Cabinet. Murali et al. [9] proposes a neural collision checker for partially observed objects and scenes. We re-implement the same pipeline as above, but use MPPI [44] iterations with Cabinet’s neural collision checking and waypoints proposals as the motion planning module.
End-to-End Behavior Cloning. We train an end-to-end policy to grasp the object from the cluster with expert demos. In our architecture, we first encoder the segmented point cloud into 1-D embeddings and concatenate them with joint states and end-effector embeddings. These embeddings are then passed to a backbone to output the delta movements. We use two variants for the backbone: an MLP [19] that takes the current step embeddings, and a transformer [45] that takes in a history of embeddings. We name these E2EImit-MLP and E2EImit-Transformer, respectively. Please refer to Appendix B.2 for details.
@@ 6.1 Baseline Comparisons
Next, we observe that the end-to-end behavior cloned policies have <<<10% success rate. Specifically, we find that the grasping phase presents significant challenges for an end-to-end trained model, as the model must learn both collision-free planning and generalize to grasp various object shapes. Thus, our dataset and benchmark offer a challenging testbed for researching imitation learning methods in object fetching, where the methods can be evaluated based on their ability to generalize to unknown objects and scenes.
@@
Table 2: Evaluation of baselines on our benchmark. We show the average success rate, the average computation time and trajectory c-space length on the success cases.
Total Succ Computation Time (s) C-Space Length (Rad)
@@
E2EImit-MLP 0.082 58.6 38.4
E2EImit-Transformer 0.099 57.5 22.3
CGN-CuRobo-Imit-MLP 0.200 17.0 20.7
CGN-CuRobo-Imit-Transformer 0.203 16.1 6.32
@@ 7 Real Robot Experiment
In total, we achieve a success rate of 21%, including 38% on on-table tasks, 13.3% on on-shelf tasks and 12.5% on in-basket tasks. The result suggests the challenge of fetching unknown objects from complex and cluttered scenes (e.g., shelves, baskets) in the real-world. Moreover, out of all the failure cases, 25.8 % comes out of low-quality grasp poses (the object slips away), 27.3 % comes out of motion planning failure, 23.5 % comes from unexpected collision with the scene, and the remaining comes from no grasp proposals. We refer to Appendix D for the experiment details, evaluations and videos.
@@ 8 Discussion and Limitation
Acknowledgments
This work was partially supported by the National Science Foundation.
@@ B.2 Imitation Learning
For our imitation learning models, we use PointNet++ [46] and Voxel-CNN [47] as the encoder on segmented point cloud of the target object, the scene and the robot. Furthermore, we encode the proprioceptive states with a MLP encoder. We concatenate all embeddings as the final embedding to feed into the policy network. For the MLP variant, we use the embedding of current step and use a 3-layer MLP to output the delta movement. The output is parameterized as a Tanh Gaussian distribution. For the Transformer variant, we use the consecutive last 4 steps’ embeddings to feed into an Optimus policy head [45] and output a Tanh Gaussian mixture distribution.
Additionally, as expert trajectories are collected in two stages, i.e., approaching and retrieval, we augment the model input with one additional one-hot vector to indicate the stage during training. Furthermore, we require the model to output one additional signal of whether the stage should come to an end (termination output). Namely, for all the training data, the very last step of each trajectory stage will have value $11$ otherwise 11-1- 1.
Though the model is trained for different stages, for E2EImit variants, we use the model in an end-to-end mode with re-grasping behavior. To be specific, if the model is in approaching stage and it outputs the termination signal, then it transits to the retrieval stage and close the gripper. However, if the gripper is fully closed during the retrieval stage (we know the grasp is a failure), then we open the gripper and transit back to the approaching stage. For the CGN-CuRobo-Imit methods, we only use it in the retrieval stage.
All models are trained on our demonstration dataset with Adam [48] optimizer. We select the best hyper-parameters by validating on held-out validation tasks. In Appendix C, we show more results on our imitation learning models.
Appendix C Additional Benchmark Experiment
We provide additional experimental results on FetchBench.
C.1 Other Manipulators
Our benchmark uses Franka Research 3 robot. However, we should note that the benchmark environments can be easily extended to other robots.
Different robot hardware will affect the overall performance of the fetching task. In Table 5, we evaluate two different hardware setups: Kinova Gen3 with Franka Hand and Kinova Gen 3 with Robotiq Hand. For simplicity, we use the ablation experiments of CGN-CuRobo.
We found that altering the arm morphology alone only slightly impacts performance due to variations in motion planning problems. However, changing the gripper significantly affects performance. This is because grasp annotations were evaluated using the Franka Hand [41], while the Robotiq gripper has a different morphology and different physical process when closing the gripper.
Table 5: Results of success rate in ablation experiments on different robot hardwares.
Franka Arm Kinova Gen3 Kinova Gen3
Franka Hand Franka Hand Robotiq Hand
GA-Mesh-Curobo 0.671 0.629 0.471
GA-Ptd-Curobo 0.190 0.173 0.108
CGN-Curobo 0.094 0.088 0.017
C.2 Imitation Learning Baselines
[…]
Table 6: Results of imitation learning baselines with different inference mode.
Success Computation Time C-Space Trajectory
E2EImit-MLP 0.082 58.59 38.42
TwoStageImit-MLP 0.053 16.54 20.88
E2EImit-Transformer 0.099 57.7 45.1
TwoStageImit-Transformer 0.065 16.67 10.30
In addition, we also evaluate the performance with twice amount of training data and with ACT [49] algorithm. The results are shown in Table 7. Training with twice amount data and on more advanced algorithms do improve the performance. We leave it as future work to improve the end-to-end imitation learning models on challenging fetching tasks.
Table 7: Results of imitation learning methods under different models and with more training data.
Success (1x Data) Success (2x Data)
TwoStageImit-MLP 0.058 0.061
E2EImit-MLP 0.082 0.093
E2EImit-ACT 0.093 -
Appendix D Real Robot Experiment
D.1 Experiment Setup
@@ D.2 Experiment Results
Appendix E Additional Related Works